# A two-minute burst of highly polarised radio emission originating from low Galactic latitude

### v1

@@ Abstract
Several sources of repeating coherent bursts of radio emission with periods of many minutes have now been reported in the literature. These “ultra-long period” (ULP) sources have no clear multi-wavelength counterparts and challenge canonical pulsar emission models, leading to debate regarding their nature. In this work we report the discovery of a bright, highly-polarised burst of radio emission at low Galactic latitude as part of a wide-field survey for transient and variable radio sources. ASKAP J175534.9--252749.1 does not appear to repeat, with only a single intense two-minute $\sim 200$ mJy burst detected from 60 hours of observations. The burst morphology and polarisation properties are comparable to those of classical pulsars but the duration is more than one hundred times longer, analogous to ULPs. No comparable bursts are detected in the rest of our widefield survey to date. Combined with the existing ULP population, this suggests that these sources have a strong Galactic latitude dependence and hints at an unexplored population of transient and variable radio sources in the thin disk of the Milky Way. The resemblance of this burst with both ULPs and pulsars calls for a unified coherent emission model for objects with spin periods from milliseconds to tens of minutes. However, whether or not these are all neutron stars or have the same underlying power source remains open for debate.
@@
pubyear: 2024pagerange: A two-minute burst of highly polarised radio emission originating from low Galactic latitudeE
@@ 1 Introduction
Searches for transient radio sources, independent of any multi-wavelength trigger, have traditionally been hindered by a combination of sensitivity and field of view. However this has changed with the advent of the Square Kilometre Array (SKA) Pathfinders – the Murchison Widefield Array (MWA; Tingay et al., 2013; Wayth et al., 2018), the Low Frequency Array (LOFAR; van Haarlem et al., 2013), MeerKAT Jonas & MeerKAT Team (2016), and the Australian SKA Pathfinder (ASKAP; Hotan et al., 2021). All-sky surveys such as the Very Large Array Sky Survey (VLASS; Lacy et al., 2020) and the Rapid ASKAP Continuum Survey (RACS; McConnell et al., 2020) have also led to the possibility of serendipitous transient searches. Millisecond timescale coherent transients such as fast radio bursts (FRBs) are now found regularly (e.g. CHIME/FRB Collaboration et al., 2021) and searches for synchrotron afterglows on timescales of weeks–years have also yielded several detections (e.g. Dong et al., 2021; Dobie et al., 2022; Dong & Hallinan, 2023).
@@
In addition, several sources of repeating coherent bursts with periods of minutes, much longer than canonical pulsar rotation periods, have now been reported. GLEAM-X J1627 has a period of 18.18 minutes (Hurley-Walker et al., 2022), GPM J1839--10 has a period of 22 minutes (Hurley-Walker et al., 2023) and ASKAP J1935+2148 has a period of of 54 minutes (Caleb et al., 2024). These three new sources are remarkably similar to the Galactic-centre radio transient GCRT J1745--3009, which has a period of 77 minutes (Hyman et al., 2005). None of the sources have detectable persistent radio emission, nor any multi-wavelength counterparts. For the purposes of this work, we consider these sources to be four members of the emerging class of ultra-long period sources (ULPs) [^1] Caleb et al. (2022) reported the discovery of a pulsar with a 76 s period, although this has been suggested to have a canonical pulsar origin (e.g. Rea et al., 2024)..
@@ 2.1 The Variables And Slow Transients (VAST) survey
In this work we use the post-processed (Jiang et al., 2024) VAST data, which has been corrected for flux scale and astrometry relative to RACS and cropped to a central $6.3\degr\times 6.3\degr$ square.
@@ 2.2 Discovery
We discovered ASKAP J1755 while searching VAST observations for sources with a single, circularly polarised, detection, as described in Appendix A. ASKAP J1755 was detected in SB47253 on 2023-01-21 with a flux density of $25.9\pm 0.8$ mJy and a circular polarisation fraction of $\sim$25%. Its position is covered by two overlapping adjacent fields, meaning that there are 85 observations of the position in our sample.
@@
Figure 2: Bottom: frequency-averaged lightcurve for ASKAP J1755, in Stokes $I$ (black), linear polarisation corrected for Faraday rotation (“$L$”; red), and Stokes $V$ (blue). Top: the linear polarisation position angle, $\Psi$. Position angle measurements are only shown where they are statistically significant. Data are taken from SB47253, which started on 2023-01-21T00:51:45.
We downloaded the calibrated measurement set of SB47253 beam 33 (the beam most sensitive at the position of the burst) from CASDA and generated dynamic spectra using DStools [^3] https://github.com/joshoewahp/dstools/. A detailed description of this process is available in Appendix B. The resulting dynamic spectra (Figure 1) show that the observed radio emission consists of a short burst with both circular and linear polarisation components. We performed Faraday rotation synthesis (Brentjens & de Bruyn, 2005a) and found a rotation measure of $\rm{RM}=961\pm 45$ rad m-2.
@@
We also fitted the Stokes $I$ dynamic spectrum with a dispersed Gaussian model and find a tentative dispersion measure of $710^{+200}_{-180}$ pc cm-3 (see Appendix C for details). Using the YMW16 model for the Milky Way electron density (Yao et al., 2017), this corresponds to a distance of $\sim 4.7$ kpc and a luminosity of $\sim 5\times 10^{21}$ erg s-1 Hz-1. The magnitudes of dispersion measure and rotation measure are consistent with Galactic plane lines of sight probed by e.g. radio pulsars (e.g. Figure S8 of Ryder et al., 2023).
Based on the inferred burst distance, brightness, and duration we estimate a brightness temperature to be $T_{b}\approx 3\times 10^{1}2$ K $D_{4.7}^{2}$, where $D=4.7D_{4.7}$ kpc is the distance to the burst source. This brightness temperature confirms the burst to have been produced by a coherent emission process. The brightness temperature, timescale, luminosity and spectral index are all comparable to the known ULP population (Hyman et al., 2005; Hurley-Walker et al., 2022, 2023; Caleb et al., 2024).
@@ 2.3.1 MeerKAT
We observed the position of the burst on 2023-05-28 with the L-band receiver on the MeerKAT radio telescope under a Director’s Discretionary Time (DDT) proposal (proposal ID: DDT-20230525-DD-01). The observing block consisted of $11\times 30$ min scans on the target, interleaved with 60 s scans of the phase calibrator J1833--2103. We used $2\times 300$ s scans of the radio galaxy J1939--6342 for bandpass and flux calibration, and a $200$ s scan of J1331+3030 for polaristion calibration. We performed standard flagging, calibration and imaging with the IDIA MeerKAT pipeline [^4] https://github.com/idia-astro/pipelines to form a deep image, with a noise level of $\sim 25\,\mu$Jy in the region around the burst. The noise is dominated by the complex extended emission associated with the Galactic Centre. We detect no emission at the transient location, with a $3\sigma$ upper limit of $75\,\mu$Jy. We formed dynamic spectra using the same process as in Section B. We find no evidence for any emission at the burst location in the 8 s/1 MHz dynamic spectrum with a $3\sigma$ upper limit of $750\,\mu$Jy.
@@ 3.2 Partially coherent mode addition
Fig. 5 provides a visual comparison between ASKAP J1755 and the output of the partial coherence model generated from these parameters. The similarity is striking. Comparing the Stokes parameters for the data and model more carefully in Fig. 6, we see that the model does not completely replicate the Stokes parameters as a fraction of total intensity: the observed polarisation fraction is slightly lower than that modelled. This may mean that there is an additional source of unpolarized radio emission in addition to that generated by partially coherent mode addition, or alternatively it may be that the mathematical implementation of the partial coherence model is not completely accounting for all of the incoherent mode addition, possibly due to the model-simplifying choice of having only one coherence fraction parameter $C$ (see appendix A2 of Oswald et al., 2023b). However, the model successfully recreates the Stokes parameters as a fraction of total polarisation: the simultaneous rotation of the PA and the presence of large amounts of circular polarisation are fully accounted for. Overall, the implication of being able to model the polarisation state of ASKAP J1755 so successfully is that it must, at least in part, have been influenced by propagation of the radio waves through a birefringent medium, which would enable orthogonal modes to be produced, and to interact.
@@
Figure 5: Left: Visualization of ASKAP J1755 527 to 617 seconds after the start of the observation. At the top is the PA profile. Below it is the total intensity (black), linear (red) and circular (blue) polarisation components, Stokes $Q$ (pale blue dotted) and Stokes $U$ (pale orange dashed). Right: a modelled representation of ASKAP J1755 using the partially coherent model with elliptical modes, for which the phase offset between the modes increases linearly over time. The model only contains information about relative mode intensity, so we scale the output by the total intensity of ASKAP J1755 (plotted in black) to generate Stokes parameters for comparison. In the PA plot, we mark the intrinsic model PA with a horizontal line at 45°$\degree$, and plot the inferred observed PA with black points.
@@ 4.1 Classification
Such a low duty cycle may be intrinsic or may be the product of the bursts being intermittent or having a flat energy distribution. Given that GPM J1839--10, GCRT J1745--3009 and ASKAP J1935+2148 are all highly intermittent, we interpret the low duty cycle as originating from intermittency. Assuming ASKAP J1755 was active for 16 days either side of the detected burst (comparable to GPM J1839--10 being active for two one-month intervals) we can conclusively rule out periods up to 43 minutes (implying a pulse duty cycle $\lesssim 5\%$), and most periods up to 2 hours (implying a pulse duty cycle $\lesssim 2\%$), based on the non-detections during that time. Motivated by GCRT J1745--3009 being active for at least 6 hours, we also consider a more conservative scenario, relying only on the observation in which it was detected and the observation of an adjacent overlapping field 27 minutes prior. In this scenario we can conclusively rule out periods up to 18.5 minutes and from 24–37 minutes (ruling out duty cycles $\gtrsim 10\%$ and from $\sim 8\%$ to $\sim 5\%$). The range of potential duty cycles in either scenario are broadly consistent with the RVM parameters measured in Section 3.1. We conclude that the data are consistent with an intermittent ULP origin, although given the lack of repeat bursts we cannot conclusively determine this.
@@
Finally, the radio detections of stellar radio bursts can usually be confidently associated with the responsible star in optical or infra-red imaging. Non-detection of an optical/infra-red counterpart therefore usually implies a distant star or cool star as the source of the emission. Taking the $J>19.8$ infra-red limit (Table 1), and the catalogue of nearby ultra-cool dwarfs from (Reid et al., 2008), we determine distance lower limits on low-mass stellar counterparts of 100similar-toabsent100$\sim$ 100∼ 100900900900900 pc for L8 to M6 spectral types. Based on a radio-to-bolometric luminosity ratio $L_{\text{rad}}/L_{\text{bol}}\sim 10^{-5}$ (Berger et al., 2010), these distances would imply radio flux densities $<90\,\mu\text{Jy}$ for a typical radio-emitting ultra-cool dwarf, which is a factor of $\sim 3000$ lower than the peak flux density of ASKAP J1755. We therefore rule out a low-mass stellar origin.
@@
4.2 Rates
We have since carried out three independent searches for sources similar to ASKAP J1755. A search of 1271 VAST-Galactic observations revealed six single-detection sources including ASKAP J1755. While some of the other five show similar properties to ASKAP J1755 (e.g. consisting of a short, polarised flare) none exhibit the distinct polarisation angle swing seen in ASKAP J1755. Notably, three of the five are located at Galactic latitudes $\lesssim 1\degr$. A detailed analysis of these sources is ongoing, but beyond the scope of this work. Expanding that search to allow $\leq 3$ detections (corresponding to a detection rate of $\sim 10\%$) to allow for periodic sources with multiple detections did not reveal any further sources. We also searched 1038 images from the VAST pilot survey (Murphy et al., 2021) for single-detection sources – after removing image artefacts we were left with no candidates. Finally we searched 1850 images from the VAST-extragalactic survey for single-detection sources, which resulted in one source associated with a known star, ASAS J055115+0355.3. Comprehensive details of all three searches are available in Appendix A.
In summary, have searched 4159 images, each covering $\sim 40\,\deg^{2}$ and found no other sources comparable to ASKAP J1755.. We found no other sources that satisfy that criteria, which is unusual given the high significance with which this burst was detected. We propose a simple, but striking, explanation – that this source originates from a currently unexplored population of transient and variable radio sources confined to the Galactic thin disk.
@@ 4.2 Rates
ASKAP J1755 is located at a Galactic latitude of $b=0.12\degr$ and the four ULP sources in the literature (Hyman et al., 2005; Hurley-Walker et al., 2022, 2023; Caleb et al., 2024) are all located within $\sim 3$ degrees of the Galactic plane. The thin disk has a scale height of 279 pc (Vieira et al., 2023), corresponding to an angular size of $\sim 2\degr$ at the Galactic Centre. The sources with distance estimates (ASKAP J1755, GLEAM-X J1627, GPM J1839--10, ASKAP J1935+2148) lie comfortably within the thin disk, while those without are likely located on the near-side of the Galactic centre and therefore also lie within the thin disk.
@@
Interpretation of the positions of the existing sample of ULPs is subject to some level of selection bias – e.g. GCRT J1745--3009 was discovered in a targeted search of the Galactic Centre and GPM J1839--10 was discovered in a search of the Galactic plane covering $|b|<15\degr$. However, the VAST survey covers $\sim 1600\,\deg^{2}$ of the Galactic plane, with coverage extending up to $|b|\sim 10\degr$, along with 265 other extragalactic fields covering $\sim 10,000\,\deg^{2}$. Hence, our search is less affected by a potential positional selection bias than previous searches. Even after accounting for a Galactic latitude dependence, it is still somewhat surprising that we have only made a single high significance detection. However this conundrum is not unprecedented – the first detection of a FRB was over 100 times the detection threshold but fainter analogues were initially elusive (Lorimer et al., 2007).
@@
Approximately 85% of the Milky Way’s stars are located in the thin disk (Anguiano et al., 2020; Allende Prieto, 2010), so a population of transients with a strong Galactic latitude dependence is not entirely unexpected – indeed, it would almost be surprising if there was not one. However, the observed distribution is so far inconsistent with the distribution of classical pulsars which are found at a wider range of Galactic latitudes due to high-velocity kicks imparted during their formation (Hobbs et al., 2005). It is also inconsistent with the observed distribution of white dwarfs, which spans a much wider range of galactic latitudes (Gentile Fusillo et al., 2021)

### v2

@@ Abstract
Several sources of repeating coherent bursts of radio emission with periods of many minutes have now been reported in the literature. These “ultra-long period” (ULP) sources have no clear multi-wavelength counterparts and challenge canonical pulsar emission models, leading to debate regarding their nature. In this work we report the discovery of a bright, highly-polarised burst of radio emission at low Galactic latitude as part of a wide-field survey for transient and variable radio sources. ASKAP J175534.9--252749.1 does not appear to repeat, with only a single intense two-minute $\sim$200-mJy burst detected from 60 hours of observations. The burst morphology and polarisation properties are comparable to those of classical pulsars but the duration is more than one hundred times longer, analogous to ULPs. Combined with the existing ULP population, this suggests that these sources have a strong Galactic latitude dependence and hints at an unexplored population of transient and variable radio sources in the thin disk of the Milky Way. The resemblance of this burst with both ULPs and pulsars calls for a unified coherent emission model for objects with spin periods from milliseconds to tens of minutes. However, whether or not these are all neutron stars or have the same underlying power source remains open for debate.
@@
pubyear: 2024pagerange: A two-minute burst of highly polarised radio emission originating from low Galactic latitudeD
@@ 1 Introduction
Searches for transient radio sources, independent of any multi-wavelength trigger, have traditionally been hindered by a combination of sensitivity and field-of-view. However this has changed with the advent of the Square Kilometre Array (SKA) Pathfinders – the Murchison Widefield Array (MWA; Tingay et al., 2013; Wayth et al., 2018), the Low Frequency Array (LOFAR; van Haarlem et al., 2013), MeerKAT Jonas & MeerKAT Team (2016), and the Australian SKA Pathfinder (ASKAP; Hotan et al., 2021). All-sky surveys such as the Very Large Array Sky Survey (VLASS; Lacy et al., 2020) and the Rapid ASKAP Continuum Survey (RACS; McConnell et al., 2020) have also led to the possibility of serendipitous transient searches. Millisecond timescale coherent transients such as fast radio bursts (FRBs) are now found regularly (e.g. CHIME/FRB Collaboration et al., 2021) and searches for synchrotron afterglows on timescales of weeks–years have also yielded several detections (e.g. Dong et al., 2021; Dobie et al., 2022; Dong & Hallinan, 2023).
@@
In addition, several sources of repeating coherent bursts with periods of minutes, much longer than canonical pulsar rotation periods, have now been reported. CHIME J0630$+$25 has a period of 7 minutes (Dong et al., 2024), GLEAM-X J1627 has a period of 18.18 minutes (Hurley-Walker et al., 2022), GPM J1839--10 has a period of 22 minutes (Hurley-Walker et al., 2023) ASKAP J1935+2148 has a period of 54 minutes (Caleb et al., 2024), ILT J1101$+$5521 has a period of $\sim 2.1$ hours (de Ruiter et al., 2024) and GLEAM-X J0704--37 has a period of $\sim$2.9 hours (Hurley-Walker et al., 2024). These six new discoveries are remarkably similar to the Galactic-centre radio transient GCRT J1745--3009, which has a period of 77 minutes (Hyman et al., 2005). None of the sources have detectable persistent radio emission. Two (ILT J1101$+$5521 and GLEAM-X J0704--37)have optical counterparts, while the others have no multi-wavelength counterparts (Kaplan et al., 2008; Rea et al., 2022, and the aforementioned discovery papers). For the purposes of this work, we consider these sources to be four members of the emerging class of ultra-long period sources (ULPs) [^1] Caleb et al. (2022) reported the discovery of a pulsar with a 76 s period, although this has been suggested to have a canonical pulsar origin (e.g. Rea et al., 2024).. The physical origin of these sources is unclear, but the leading prospects are white dwarfs or neutron stars (e.g. Tong, 2023). A number of scenarios have been proposed to explain the unusually long periods in these scenarios (e.g., wind-dominated spin-down of a regular pulsar; Sun et al., 2024), while more exotic origins have also been proposed (e.g. Baumgarte & Shapiro, 2024).
@@ 2.1 The Variables And Slow Transients (VAST) survey
In this work we use the post-processed (Jiang et al., 2024) VAST data, which has been corrected for flux scale and astrometry relative to RACS and cropped to a central $6.3\degr\times 6.3\degr$ square. We regularly ingest the post-processed data covering the period from into the VAST pipeline (Pintaldi et al., 2022) as it becomes available. The VAST pipeline takes images, noise maps and source catalogues from ASKAP observations, runs source association, carries out forced photometry where appropriate, and calculates source statistics. The resulting data can then be queried programmatically using vast-tools, a python package designed for interacting with VAST data (Stewart et al., 2023).
@@ 2.2 Discovery
As part of VAST pipeline testing we ran a search for sources with a single detection on an early subset of the VAST Galactic survey including observations through to early 2023. We selected sources that satisfied the following criteria:
  • nearest neighbour distance larger than 1 arcminute;
  • number of relations equal to zero (i.e. the source is isolated).
We then searched for sources with a single detection and then used vast-tools to determine whether that detection was circularly polarised. ASKAP J1755 was the only source that satisfied all criteria in the original search, although several other sources have been detected in more recent searches, which will be presented in future work.
ASKAP J1755 was detected in SB47253 on 2023-01-21 with a flux density of $25.9\pm 0.8$ mJy and a circular polarisation fraction of $\sim$25%. After applying astrometric corrections (see Appendix C for details), we measured a position of 17:55:34.9(1) -25:27:49.1(1), with uncertainties in the final digit quoted in brackets. This position is covered by two overlapping adjacent fields, meaning that there are 85 observations of the position in our sample.
@@
Figure 2: Bottom: frequency-averaged lightcurve for ASKAP J1755, in Stokes $I$ (black), de-biased linear polarisation corrected for Faraday rotation (“$L$”; red), and Stokes $V$ (blue). Top: the linear polarisation position angle, $\Psi$. Position angle measurements are only shown where they are statistically significant. Data are taken from SB47253, which started on 2023-01-21T00:51:45.
We downloaded the calibrated measurement set of SB47253 beam 33 (the beam most sensitive at the position of the burst) from CASDA and generated dynamic spectra using DStools (Pritchard, 2024). A detailed description of this process is available in Appendix A. The resulting dynamic spectra (Figure 1) show that the observed radio emission consists of a short burst with both circular and linear polarisation components. We performed Faraday rotation measure (RM) synthesis (Brentjens & de Bruyn, 2005a) and found an RM of $961\pm 45$ rad m-2.
@@
We also fitted the Stokes $I$ dynamic spectrum with a dispersed Gaussian model and find a tentative dispersion measure of $710^{+200}_{-180}$ pc cm-3 (see Appendix B for details), although we note that the inferred dispersion delay across our observing band is significantly less than the integration time. Using the YMW16 model for the Milky Way electron density (Yao et al., 2017), this corresponds to a distance of $\sim 4.7$ kpc and a luminosity of $\sim 5\times 10^{21}$ erg s-1 Hz-1. The magnitudes of dispersion measure and rotation measure are consistent with Galactic plane lines of sight probed by e.g. radio pulsars (e.g. Figure S8 of Ryder et al., 2023).
Based on the inferred burst distance, brightness, and duration we estimate a brightness temperature to be $T_{b}\approx 3\times 10^{12}$ K $D_{4.7}^{2}$, where $D=4.7D_{4.7}$ kpc is the distance to the burst source. This brightness temperature confirms the burst to have been produced by a coherent emission process. The brightness temperature, timescale, luminosity and spectral index are all comparable to the known ULP population (Hyman et al., 2005; Hurley-Walker et al., 2022, 2023; Caleb et al., 2024; de Ruiter et al., 2024; Hurley-Walker et al., 2024; Dong et al., 2024).
@@ 2.3.1 MeerKAT
We observed the position of the burst on 2023-05-28 with the L-band receiver on the MeerKAT radio telescope under a Director’s Discretionary Time (DDT) proposal (proposal ID: DDT-20230525-DD-01). The observing block consisted of $11\times 30$ min scans on the target, interleaved with 60 s scans of the phase calibrator J1833--2103. We used $2\times 300$ s scans of the radio galaxy J1939--6342 for bandpass and flux calibration, and a $200$ s scan of J1331+3030 for polaristion calibration. We performed standard flagging, calibration and imaging with the IDIA MeerKAT pipeline [^3] https://github.com/idia-astro/pipelines to form a deep image, with a noise level of $\sim$25 $\mu$Jy/beam in the region around the burst. The noise is dominated by the complex extended emission associated with the Galactic Centre. We detect no emission at the transient location, with a $3\sigma$ upper limit of $75\,\mu$Jy. We formed dynamic spectra using the same process as in Section A. We find no evidence for any emission at the burst location in the 8 s/1 MHz dynamic spectrum with a $3\sigma$ upper limit of $750\,\mu$Jy.
@@ 3.2 Partially coherent mode addition
Fig. 5 provides a visual comparison between ASKAP J1755 and the output of the partial coherence model generated from these parameters. Comparing the Stokes parameters for the data and model more carefully in Fig. 6, we see that the model does not completely replicate the Stokes parameters as a fraction of total intensity: the observed polarisation fraction is slightly lower than that modelled. This may mean that there is an additional source of unpolarized radio emission in addition to that generated by partially coherent mode addition, or alternatively it may be that the mathematical implementation of the partial coherence model is not completely accounting for all of the incoherent mode addition, possibly due to the model-simplifying choice of having only one coherence fraction parameter $C$ (see appendix A2 of Oswald et al., 2023b). However, the model successfully recreates the Stokes parameters as a fraction of total polarisation: the simultaneous rotation of the PA and the presence of large amounts of circular polarisation are fully accounted for. Overall, the implication of being able to model the polarisation state of ASKAP J1755 so successfully is that it must, at least in part, have been influenced by propagation of the radio waves through a birefringent medium, which would enable orthogonal modes to be produced, and to interact.
@@
Figure 5: Polarised radio emission from ASKAP J1755 around the time of the burst. The top panel shows the polarisation angle profile and the bottom panel shows the light curve for Stokes $I$ (black), $Q$ (orange), $U$ (red) and $V$ (blue). The markers in both panels show the observed data, while the lines show the model as described in Section 3.2. Uncertainties on the observed data are included, but in most cases are smaller than the marker size. The model only contains information about relative mode intensity, so we scale the output by the total intensity of ASKAP J1755 to generate Stokes parameters for comparison. In the PA plot, we mark the intrinsic model PA with a horizontal dashed line at 45°$\degree$.
@@ 4.1 Classification
Such a low duty cycle may be intrinsic or may be the product of the bursts being intermittent or having a flat energy distribution. Given that GPM J1839--10, GCRT J1745--3009 and ASKAP J1935+2148 are all highly intermittent, we interpret the low duty cycle as originating from intermittency. Assuming ASKAP J1755 was active for 16 days either side of the detected burst (comparable to GLEAM-X J1627 being active for two one-month intervals, and motivated by the available VAST observations) we can conclusively rule out periods up to 43 minutes (implying a pulse duty cycle $\lesssim 5\%$), and most periods up to 2 hours (implying a pulse duty cycle $\lesssim 2\%$), based on the non-detections during that time. Motivated by GCRT J1745--3009 being active for at least 6 hours, we also consider a more conservative scenario, relying only on the observation in which it was detected and the observation of an adjacent overlapping field 27 minutes prior. In this scenario we can conclusively rule out periods up to 18.5 minutes and from 24–37 minutes (ruling out duty cycles $\gtrsim 10\%$ and from $\sim 8\%$ to $\sim 5\%$). Details of this analysis can be found in Appendix D. The range of potential duty cycles in either scenario are broadly consistent with the RVM parameters measured in Section 3.1. We conclude that the data are consistent with an intermittent ULP origin, although given the lack of repeat bursts we cannot conclusively determine this.
@@
Finally, the radio detections of stellar radio bursts can usually be associated with the responsible star in optical or infra-red imaging. Non-detection of an optical/infra-red counterpart therefore usually implies a distant star or cool star as the source of the emission. Taking the $J>19.8$ infra-red limit (Table 1), and the catalogue of nearby ultra-cool dwarfs from (Reid et al., 2008), we determine distance lower limits on low-mass stellar counterparts of 100similar-toabsent100$\sim$ 100∼ 100900900900900 pc for L8 to M6 spectral types. Based on a radio-to-bolometric luminosity ratio $L_{\text{rad}}/L_{\text{bol}}\sim 10^{-5}$ (Berger et al., 2010), these distances would imply radio flux densities $<90\,\mu\text{Jy}$ for a typical radio-emitting ultra-cool dwarf, which is a factor of $\sim 3000$ lower than the peak flux density of ASKAP J1755. We therefore rule out a low-mass stellar origin.
@@
4.2 A population confined to the Galactic Plane
@@
ASKAP J1755 is located at a Galactic latitude of $b=0.12\degr$ and the five of the seven ULP sources in the literature (Hyman et al., 2005; Hurley-Walker et al., 2022, 2023; Caleb et al., 2024) are all located within $\sim 3$ degrees of the Galactic plane. The thin disk has a scale height of 279 pc (Vieira et al., 2023), corresponding to an angular size of $\sim 2\degr$ at the Galactic Centre. Six of the seven ULP sources (all but GCRT J1745--3009) have distance estimates, along with ASKAP J1755 and hence their position within the Galaxy can be computed. ASKAP J1755, GLEAM-X J1627, GPM J1839--10, ASKAP J1935+2148 and CHIME J0630$+$25 all lie comfortably within the thin disk while ILT J1101$+$5521 and GLEAM-X J0704--37 are located at heights of $|z|\approx 430$ pc and $|z|\approx 330\,$pc above the Galactic plane respectively. While GCRT J1745--3009 does not have a distance constraint, its extremely low galactic latitude ($|b|\sim 0.5\degr$) suggests it also lies within the thin disk.
@@
Interpretation of the positions of the existing sample of ULPs is subject to some level of selection bias – e.g. GCRT J1745--3009 was discovered in a targeted search of the Galactic Centre and GPM J1839--10 was discovered in a search of the Galactic plane covering $|b|<15\degr$. However, the VAST survey covers $\sim 1600\,\deg^{2}$ of the Galactic plane, with coverage extending up to $|b|\sim 10\degr$, along with 265 other extragalactic fields covering $\sim 10,000\,\deg^{2}$. Hence, our search is less affected by a potential positional selection bias than previous searches.
@@
Approximately 85% of the Milky Way’s stars are located in the thin disk (Anguiano et al., 2020; Allende Prieto, 2010), so a population of transients confined to the Galactic plane is not entirely unexpected – indeed, it would almost be surprising if there was not one. However, the observed distribution is so far inconsistent with the distribution of classical pulsars which are found at a wider range of Galactic latitudes due to high-velocity kicks imparted during their formation (Hobbs et al., 2005). It is also inconsistent with the observed distribution of white dwarfs, which spans a much wider range of galactic latitudes (Gentile Fusillo et al., 2021). In contrast, the observed magnetar population (which is presumably biased towards young, active objects) is found almost exclusively at very low Galactic latitudes, with a typical scale height of 20–31 pc (Olausen & Kaspi, 2014).